# Using the Empirical Attainment Function for Analyzing Single-objective Black-box Optimization Algorithms

### v1

@@ I Introduction
In the context of benchmarking sampling-based single-objective optimization algorithms, one of the most common ways to illustrate the performance on a single problem instance is to plot the objective value of a best-so-far solution as a function of the number of solutions that have been evaluated. There are different options for the aggregation of such convergence curves over different problem instances. In evolutionary computation, one of the most popular ways to aggregate performance data over several problems is based on the so-called empirical cumulative distribution function (ECDF) of the number of targets attained at a given runtime. This alternative was notably promoted by the COCO platform [1]. To define an ECDF curve, one fixes a set of target values and counts, as a function of the solutions evaluated, how many of these targets have been achieved by the algorithm. These numbers are then normalized across all runs of the algorithm on each problem instance and then across all problem instances. This target-based ECDF generalizes the concept of time-to-target [2] or runtime distribution of a Las Vegas algorithm for a decision problem [3] to multiple targets. The area under the target-based ECDF curve (AUC) can be seen as a measure of the anytime performance of the algorithm [4]. [^1] The COCO platform [1] minimizes area over the ECDF curve, which is equivalent to the maximum runtime minus the AUC, because the former approximates the geometric average runtime.
@@ II-A Target-based Empirical Cumulative Distribution Function
$\widehat{F}_{Z}(t)=\frac{1}{r}\sum_{i=1}^{r}\frac{1}{|Z|}\sum_{z\in Z}\text{1}\left(V(A,f,t,i)\leq z\right),$ (1)
@@ II-B Empirical Attainment Function
In the context of benchmarking single-objective optimizers, given a pair $(t,z)$ of budget and function value, respectively, we say that run $i$ of algorithm $A$ attains $(t,z)$ if and only if algorithm $A$ solving problem $f$ obtains an objective function value not worse than $z$ not later than $t$, that is, if and only if $V(A,f,t,i)\leq z$. We can formulate the EAF in this context as the fraction of $r$ runs of algorithm $A$ on function $f$ that attain an objective value $z$ at budget $t$:
$\widehat{\alpha}(t,z)=\frac{1}{r}\sum_{i=1}^{r}\text{1}\left(V(A,f,t,i)\leq z\right)$ (2)
The above matches the original formulation of the EAF [5], if we consider that each run $i$ produces a set $S_{i}$ of points $\left(t,V(A,f,t,i)\right)$, i.e., $S_{i}=\{\left(t,V(A,f,t,i)\right)\mid 1\leq t\leq t_{\text{max}}\}$. In this original formulation, the value of the EAF of a set $\{S_{1},\dotsc,S_{n}\}$ of sets $S_{i}\subset\mathbb{R}^{d}$ at a given point $y=(y_{1},\dotsc,y_{d})\in\mathbb{R}^{d}$ is defined as the fraction of sets $S_{i}$ for which there exists at least one point $x=(x_{1},\dotsc,x_{d})\in S_{i}$ such that, for all $1\leq j\leq d$ it holds that $x_{j}\leq y_{j}$. In this case, we say that set $S_{i}$ attains the point $y$.
See Figure 2 for an illustrative example of the EAF for three sets, corresponding to three runs of an algorithm.
@@ II-D Area under the EAF as a measure of anytime performance
$\text{AUC}\left(F_{\widehat{\alpha}}(t)\right)=\frac{1}{z_{\text{max}}-z_{\text{min}}}\int_{1}^{t_{\text{max}}}\mathop{}\!\mathrm{d}t\int_{z_{\text{min}}}^{z_{\text{max}}}\widehat{\alpha}(t,z)\mathop{}\!\mathrm{d}z$ (7)
@@
$\text{AUC}\left(\widehat{\alpha}(t,z)\right)=\int_{(1,-\infty)}^{(t_{\text{max}},z_{\text{max}})}\widehat{\alpha}(t,z)\mathop{}\!\mathrm{d}t\mathop{}\!\mathrm{d}z$ (8)
@@
Another measure of anytime performance proposed in the literature [18] is the area over the convergence curve $S_{i}=\{(t,V(A,f,t,i))\mid 1\leq t\leq t_{\text{max}}\}$, which can be calculated by summing up the areas of the rectangles formed by the points $(t,V(A,f,t,i))$ and $(t+1,z_{\text{max}})$ as follows:
$\text{AOCC}(S_{i})=\sum_{t=1}^{t_{\text{max}}-1}\max\{0,z_{\text{max}}-V(A,f,t,i)\}$ (9)
@@
$\begin{split}\int_{1}^{t_{\text{max}}}\mathop{}\!\mathrm{d}t\int_{-\infty}^{z_{\text{max}}}\widehat{\alpha}(t,z)\mathop{}\!\mathrm{d}z\\ =\int_{1}^{t_{\text{max}}}\mathop{}\!\mathrm{d}t\int_{-\infty}^{z_{\text{max}}}\frac{1}{r}\sum_{i=1}^{r}\text{1}\left(V(A,f,t,i)\leq z\right)\mathop{}\!\mathrm{d}z\\ =\frac{1}{r}\sum_{i=1}^{r}\int_{1}^{t_{\text{max}}}\mathop{}\!\mathrm{d}t\int_{-\infty}^{z_{\text{max}}}\text{1}\left(V(A,f,t,i)\leq z\right)\mathop{}\!\mathrm{d}z\end{split}$ (10)
@@
$\begin{split}\frac{1}{r}\sum_{i=1}^{r}\int_{1}^{t_{\text{max}}}\mathop{}\!\mathrm{d}t\int_{-\infty}^{z_{\text{max}}}\text{1}\left(V(A,f,t,i)\leq z\right)\mathop{}\!\mathrm{d}z\\ =\frac{1}{r}\sum_{i=1}^{r}\sum_{t=1}^{t_{\text{max}}-1}\max\{0,z_{\text{max}}-V(A,f,t,i)\}\\ =\frac{1}{r}\sum_{i=1}^{r}\text{AOCC}(S_{i})\end{split}$ (11)
In other words, not only the AUC of the EAF-based ECDF is equivalent to the AUC of the EAF (multiplied by a normalization constant), but also both are equivalent to the mean AOCC of the runs used to compute the EAF. There are situations in which we are only interested in the AUC value, for example, for benchmarking purposes or when using the AUC as the metric that guides the tuning of parameters. In those situations, instead of storing the data for all runs to compute the target-based ECDF or the EAF and their AUC value, it is much more convenient to compute the AOCC of each run as soon as it finishes and only store the AOCC values. Moreover, recomputing the mean AOCC after performing additional runs is trivial, whereas updating the EAF or the ECDF is less so. The use of the mean AOCC to guide the tuning of parameters was already demonstrated in the literature [18], but the connection with the AUC of the target-based ECDF was not known at the time.
@@ III-B Integration into IOHanalyzer
For practical usage of the EAF with various types of benchmarking data, we have integrated several ways of using EAF-based analysis into IOHanalyzer [11], which is a part of the IOHprofiler benchmarking platform. For this purpose, the computations of the EAF are calculated by version 2.4.1 of the eaf R package [7]. [^2] https://mlopez-ibanez.github.io/eaf/ Through this integration, several ways of computing, visualizing and comparing EAF-based statistics, are available directly in the web-interface of IOHananlyzer (and the corresponding R package). As such, benchmarking data from platforms such as COCO [1], Nevergrad [20], and IOHexperimenter [21] can also be analyzed using these methods.
To speed up computations when creating visualizations, we make use of a subsampling approach when calculating the EAF, ECDF, and AUC values. This subsampling is done on the runtime-axis, where we select a set of $50$ log-spaced runtimes at which we calculate the fraction of targets hit for the target-based ECDF or at which we take the function values to pass to the EAF calculation.
@@ IV Use in Practice
In addition to the individual EAF plots, we can also perform a comparative analysis by plotting the differences between two EAFs. This is illustrated in Figure 5, where the differences between CMA-ES and BFGS on a single BBOB function are shown. In this example, BFGS outperforms CMA-ES between $50$ and ${10}\mathrm{e}{3.5}$ evaluations, whereas CMA-ES wins after that, managing to converge to $f(x)=10^{-8}$ more reliably. This analysis can additionally be extended to highlight the difference between an algorithm and a portfolio, by taking the upper envelope of the portfolio’s EAF. This portfolio-based comparison is illustrated in Figure 6, where a set of four algorithms are compared on benchmark function F11. The selected portfolio consists of CMA-ES, BFGS, Differential Evolution [29] (DE) and Cobyla [30], and their performance data is taken from the COCO data repository. From this figure, we see that in three stages of the search, Cobyla, BFGS, and CMA-ES, respectively, outperform every other considered algorithm, whereas there is no pair $(t,z)$ with $t>10$ in which the EAF of DE dominates the maximum of the EAF values of the other three algorithms.
@@ IV-A Sensitivity of the ECDF wrt Target Values
Figure 7: ECDF as calculated using the EAF (4) and the target-based approach (1), with different numbers of targets, for two different algorithms, DE and MLSL. Data is aggregated over all 24 10-dimensional BBOB functions.
@@
In Figure 7, we illustrate the impact of changing the number of targets for two of these algorithms, DE and Multi-Level Single-Linkage (MLSL) [31]. As can be seen from this figure, the ECDF based on a limited number of targets consistently overestimates the EAF-based ECDF.
@@ VI Conclusions
Acknowledgment
This work was initiated when M. López-Ibáñez held an invited Professorship at Sorbonne University, France. Part of this work was carried out during Dagstuhl Seminar 23251 “Challenges in Benchmarking Optimization Heuristics” and at the “Benchmarked: Optimization Meets Machine Learning” workshop at the Lorentz center.

### v2

@@ I Introduction
In the context of benchmarking sampling-based single-objective optimization algorithms, one of the most common ways to illustrate the performance on a single problem instance is to plot the objective value of a best-so-far solution as a function of the number of solutions that have been evaluated. There are different options for the aggregation of such convergence curves over different problem instances. In evolutionary computation, one of the most popular ways to aggregate performance data over several problems is based on the so-called empirical cumulative distribution function (ECDF) of the number of targets attained at a given runtime. This alternative was notably promoted by the COCO platform [1]. To define an ECDF curve, one fixes a set of target values and counts, as a function of the solutions evaluated, how many of these targets have been achieved by the algorithm. These numbers are then normalized across all runs of the algorithm on each problem instance and then across all problem instances. The area under the target-based ECDF curve (AUC) can be seen as a measure of the anytime performance of the algorithm [2]. [^1] The COCO platform [1] minimizes area over the ECDF curve, which is equivalent to the maximum runtime minus the AUC, because the former approximates the geometric average runtime.
This target-based ECDF generalizes the concept of time-to-target [3] or runtime distribution of a Las Vegas algorithm for a decision problem [4] to multiple targets.
@@ II-A Target-based Empirical Cumulative Distribution Function
$\widehat{F}_{Z}(t)=\frac{1}{r}\sum_{i=1}^{r}\frac{1}{|Z|}\sum_{z\in Z}\text{1}\left(V(A,f,t,i)\leq z\right)\enspace,$ (1)
@@ II-B Empirical Attainment Function
In the context of benchmarking the anytime performance of single-objective optimizers, given a pair $(t,z)$ of budget and function value, respectively, we say that run $i$ of algorithm $A$ attains $(t,z)$ if and only if algorithm $A$ solving problem $f$ obtains an objective function value not worse than $z$ not later than $t$, that is, if and only if $V(A,f,t,i)\leq z$. We can formulate the EAF in this context as the fraction of $r$ runs of algorithm $A$ on function $f$ that attain an objective value $z$ at budget $t$:
$\widehat{\alpha}(t,z)=\frac{1}{r}\sum_{i=1}^{r}\text{1}\left(V(A,f,t,i)\leq z\right)\enspace.$ (2)
The above matches the original formulation of the EAF [5], if we consider that each run $i$ produces a set $S_{i}\subset\mathbb{R}^{2}$ of points $\left(t,V(A,f,t,i)\right)$, i.e., $S_{i}=\{\left(t,V(A,f,t,i)\right)\mid 1\leq t\leq t_{\text{max}}\}$. In this original formulation, a set $S_{i}\subset\mathbb{R}^{d}$ attains the point $y=(y_{1},\dotsc,y_{d})\in\mathbb{R}^{d}$ iff there exists at least one point $s=(s_{1},\dotsc,s_{d})\in S_{i}$ such that, for all $1\leq j\leq d$, it holds that $s_{j}\leq y_{j}$. In this case, the value of the EAF of a set of sets $\{S_{1},\dotsc,S_{n}\}$ at a given point $y\in\mathbb{R}^{d}$ is defined as the fraction of sets $S_{i}$, $i=1,\dotsc,n$, that attain that point $y$. See Figure 2 for an illustrative example of the EAF for three sets, corresponding to three runs of an algorithm.
@@ II-D Area under the EAF as a measure of anytime performance
$\text{AUC}\left(F_{\widehat{\alpha}}(t)\right)=\frac{1}{z_{\text{max}}-z_{\text{min}}}\int_{1}^{t_{\text{max}}}\mathop{}\!\mathrm{d}t\int_{z_{\text{min}}}^{z_{\text{max}}}\widehat{\alpha}(t,z)\mathop{}\!\mathrm{d}z\enspace.$ (7)
@@
$\text{AUC}\left(\widehat{\alpha}(t,z)\right)=\int_{(1,-\infty)}^{(t_{\text{max}},z_{\text{max}})}\widehat{\alpha}(t,z)\mathop{}\!\mathrm{d}t\mathop{}\!\mathrm{d}z\enspace,$ (8)
@@
Another measure of anytime performance proposed in the literature [18] is the area over the convergence curve (AOCC), $S_{i}=\{(t,V(A,f,t,i))\mid 1\leq t\leq t_{\text{max}}\}$, which can be calculated by summing up the areas of the rectangles formed by the points $(t,V(A,f,t,i))$ and $(t+1,z_{\text{max}})$ as follows:
$\text{AOCC}(S_{i})=\sum_{t=1}^{t_{\text{max}}-1}\max\{0,z_{\text{max}}-V(A,f,t,i)\}\enspace.$ (9)
@@
$\begin{split}\int_{1}^{t_{\text{max}}}\mathop{}\!\mathrm{d}t\int_{-\infty}^{z_{\text{max}}}\widehat{\alpha}(t,z)\mathop{}\!\mathrm{d}z\\ =\int_{1}^{t_{\text{max}}}\mathop{}\!\mathrm{d}t\int_{-\infty}^{z_{\text{max}}}\frac{1}{r}\sum_{i=1}^{r}\text{1}\left(V(A,f,t,i)\leq z\right)\mathop{}\!\mathrm{d}z\\ =\frac{1}{r}\sum_{i=1}^{r}\int_{1}^{t_{\text{max}}}\mathop{}\!\mathrm{d}t\int_{-\infty}^{z_{\text{max}}}\text{1}\left(V(A,f,t,i)\leq z\right)\mathop{}\!\mathrm{d}z\enspace.\end{split}$ (10)
@@
$\begin{split}\frac{1}{r}\sum_{i=1}^{r}\int_{1}^{t_{\text{max}}}\mathop{}\!\mathrm{d}t\int_{-\infty}^{z_{\text{max}}}\text{1}\left(V(A,f,t,i)\leq z\right)\mathop{}\!\mathrm{d}z\\ =\frac{1}{r}\sum_{i=1}^{r}\sum_{t=1}^{t_{\text{max}}-1}\max\{0,z_{\text{max}}-V(A,f,t,i)\}\\ =\frac{1}{r}\sum_{i=1}^{r}\text{AOCC}(S_{i})\enspace.\end{split}$ (11)
In other words, not only the AUC of the EAF-based ECDF is equivalent to the AUC of the EAF (multiplied by a normalization constant), but also both are equivalent to the mean AOCC of the runs used to compute the EAF. A more general proof of the relationship between the expected value of a generalized hypervolume indicator (i.e., where the AOCC is multiplied by a weight function) and the first-order attainment function (i.e., the EAF when the number of runs goes to infinity) is provided by [19]. The simpler proof in Eq. (10)–(11) does not consider the weighted variant of the hypervolume indicator, but it applies to a finite number of empirical samples.
There are situations in which we are only interested in the AUC value, for example, for benchmarking purposes or when using the AUC as the metric that guides the tuning of parameters. In those situations, instead of storing the data for all runs to compute the target-based ECDF or the EAF and their AUC value, it is much more convenient to compute the AOCC of each run as soon as it finishes and only store the AOCC values. Moreover, recomputing the mean AOCC after performing additional runs is trivial, whereas updating the EAF or the ECDF is less so. The use of the mean AOCC to guide the tuning of parameters was already demonstrated in the literature [18], but the connection with the AUC of the target-based ECDF was not known at the time.
@@ III-B Integration into IOHanalyzer
For practical usage of the EAF with various types of benchmarking data, we have integrated several ways of using EAF-based analysis into IOHanalyzer [11], which is a part of the IOHprofiler benchmarking platform. For this purpose, the computations of the EAF are calculated by version 2.4.1 of the eaf R package [7]. [^2] https://mlopez-ibanez.github.io/eaf/ Through this integration, several ways of computing, visualizing and comparing EAF-based statistics, are available directly in the web-interface of IOHanalyzer (and the corresponding R package). Benchmarking data from platforms such as COCO [1], Nevergrad [21], and IOHexperimenter [22] can also be analyzed using these methods.
To speed up computations when creating visualizations, we make use of a subsampling approach when calculating the EAF, ECDF, and AUC values. This subsampling is done on the runtime-axis, where we select a set of $50$ log-spaced runtimes at which we calculate the fraction of targets hit for the target-based ECDF or at which we take the function values to pass to the EAF calculation. Given this subsampling, the computation of the EAF and the EAF-based ECDF took just 432 seconds in a single CPU for all datasets considered in the next section (1295 datasets in total), which is a reasonable time for the analysis of an experiment of this scale.
@@ IV Use in Practice
In addition to the individual EAF plots, we can also perform a comparative analysis by plotting the differences between two EAFs. This is illustrated in Figure 5, where the differences between CMA-ES and BFGS on a single BBOB function are shown. In this example, BFGS outperforms CMA-ES between $50$ and $1\text{e}{+}3.5$ (approx. $3162$) evaluations, whereas CMA-ES wins after that, managing to converge to $f(x)=10^{-8}$ more reliably. This analysis can additionally be extended to highlight the difference between an algorithm and a portfolio, by taking the upper envelope of the portfolio’s EAF. This portfolio-based comparison is illustrated in Figure 6, where a set of four algorithms are compared on benchmark function F11. The selected portfolio consists of CMA-ES, BFGS, Differential Evolution [30] (DE) and Cobyla [31], and their performance data is taken from the COCO data repository. From this figure, we see that in three stages of the search, Cobyla, BFGS, and CMA-ES, respectively, outperform every other considered algorithm, whereas there is no pair $(t,z)$ with $t>10$ in which the EAF of DE dominates the maximum of the EAF values of the other three algorithms.
@@ IV-A Sensitivity of the ECDF wrt Target Values
Figure 7: ECDF as calculated using the EAF (4) and the target-based approach (1), with different numbers of targets, for two different algorithms, BFGS and CMA-ES. Data is aggregated over all 24 10-dimensional BBOB functions.
@@
In Figure 7, we illustrate the impact of changing the number of targets for the BFGS and CMA-ES. As can be seen from this figure, the ECDF based on a limited number of targets consistently overestimates the EAF-based ECDF.
@@ V EAF-based Statistics
Statistical measures and visualization techniques [35] that can be applied to the target-based ECDF are also applicable to the EAF-based ECDF. Some techniques may be directly applicable to the EAF. For example, it is possible to compute bootstrapped confidence bands [35] around the EAF that estimate the uncertainty. However, applying other techniques directly to the EAF may require additional research effort to handle the challenge of samples being sets of (runtime, quality) points rather than scalar numerical values.